# CircleZ : Reliable Photometric redshifts for AGN computed using only photometry from Legacy Survey Imaging for DESI

### v1

@@ 3.2 Extension of the training sample to high-z sources
The contribution to the high-redshift AGN population at $z$¿5 from X-ray pencil-beam surveys is feeble  [^14] Unsurprisingly, most of the sources at redshift higher than five that we visually inspected were wrong. By construction, the training sample would not represent the high-$z$ population. However, eROSITA is expected to detect about 90 sources at $z$¿5.5 by the end of the mission (Wolf et al., 2021). At the same time, about a dozen have already been identified (Medvedev et al., 2020, 2021; Wolf et al., 2021, 2023, Wolf et al., in prep). The optical spectra of these high-z QSOs closely resemble the spectra of QSOs at low redshift (e.g., Mortlock et al., 2011; Yang et al., 2021), which is why we have extended the training sample, adding 283 sources spectroscopically confirmed at $z>5.5$, from the compilation of Fan et al. (2023), regardless of whether or not they are detected in the X-rays. The violet dashed histogram in Figure 3 shows the redshift distribution of these sources.
Table 2: Breakdown distribution of sources used for training (i.e., training, test, and validation samples) and in the blind samples, before and after the match with the spectroscopic redshift compilation, and after cleaning for duplicates. The numbers within the parenthesis show the contribution to the sample used for training.
@@ 3.3 Samples for blind tests
CSC2: this blind sample is generated using the counterparts to the CSC2 objects https://cxc.cfa.harvard.edu/csc/ in the same way as described for the training sample. The original sample of CSC2 sources with reliable counterparts included 5643 sources. Screening for sources with spectroscopic redshift and after removing the sources that are in common with the training sample, 413 remain in the area covered by LS10-South, after including 26 redshifts from SDSS-V.
@@
Figure 4: X-rays flux vs. W1 magnitude distributions for the training (left panel), and the blind samples: CSC2 (middle panel) and Chandra-COSMOS (right panel). The dashed line represents the galactic/extragalactic separation proposed in (Salvato et al., 2018b), who already pointed out that bright nearby galaxies would be misclassified. The contours encompassing 50%, 80%, and 95% of the training sample distribution are overplotted on all panels. While the CSC2 sample has a distribution similar to the training sample, the distribution of the Chandra-COSMOS Legacy sources is biased toward fainter values of X-rays and W1.
@@ 4.1 Problem Formulation
However, as shown in D’Isanto et al. (2018), a single trained network producing a PDF cannot capture the uncertainty. We, therefore, extended our approach for uncertainty estimation and followed Lakshminarayanan et al. (2017) who proposed to use an ensemble of deep neural networks such as MDNs to improve uncertainty estimates. We follow this approach with an ensemble of MDNs that, given the photometric input features, each predict the parameters of the Gaussian mixture distribution with a certain number of components. To be specific, the $i^{th}$ model $f_{i}(x;\theta_{i})$ in an ensemble of $N$ models, given its parameters $\theta_{i}$ and the input features $x$, predicts the parameters of a Gaussian mixture composed of $n_{i}$ Gaussian, defined by the means $\mu^{i}(x;\theta_{i})$, and standard deviations $\sigma^{i}(x;\theta_{i})$ and weights $\omega^{i}(x;\theta_{i})$. The conditional predictive distribution defined by this MDN is given by:
@@ 4.2.2 Training procedure and hyperparameter search
Table 3: Model and training hyperparameters with the range within which they are selected. These specific ranges are selected by observing model performance with hyperparameter values in wider ranges.
@@ 4.2.3 Inference and post-processing
Table 4: Averaged statistics across all individually trained NLL loss and CPRS models. The table includes aggregated information for the combined ensemble of all NLL and CRPS models. Furthermore, the fraction of outliers and accuracy are listed for the best-performing ensembles.
@@ 5 Redshift computation for eFEDS sources and comparison with literature
It is also worth mentioning that the metrics reported for LePHARE and DNNZ were measured using the training sample, not on the test and/or validation sample. If we compute the metrics for CircleZ using all the spectroscopy, i.e., including those used for the training and validation, the values would suggest a quality close to the one available for the best pencil-beam surveys (XMM-COSMOS, (E)CDF Salvato et al., 2009; Luo et al., 2010; Hsu et al., 2014), computed with deep, rich and homogenised photometry). This assures that the new photo-zs made available for eFEDS in this paper are reliable and can be used for Luminosity Function studies or for preselecting specific instruments or wavelengths of interest when submitting a proposal.
@@
Table 5: Comparison of bias, fraction of outliers, and accuracy between photo-z computed with CircleZ, LePHARE, and DNNZ, for extragalactic sources in the eROSITA/eFEDS field. The values for LePHARE and DNNZ are taken directly from (Salvato et al., 2022) for comparison. While the selection of the templates for LePHARE is based on the same sample used for assessing the quality of photo-zs, the training for DNNz(CircleZ) is partially(completely) independent from the test and validation sample used here. The quality of CircleZ exceeds the one obtained by both LePHARE and DNNZ in the area outside KiDS, demonstrating the strength of the method in general and particularly where there is a lack of near-infrared photometry.
@@
5.1 New photo-z catalog release for eFEDS
With this paper, we release a new catalog of photo-z computed with CircleZ, for the extragalactic sources in the eROSITA/eFEDS field. For each source we provide, in addition to the eROSITA_ID, the photo-z, and the 1 and 3 $\sigma$ lower and upper errors, also a unique identifier for the counterpart in the LS10 catalog. The identifier is created by concatenating the columns RELEASE, BRICKID, and OBJID. In addition, we provide about 5000 new spectroscopic redshifts that became available with the 18th Data Release of SDSS-V (Almeida et al., 2023). The catalog is available via Vizier and on the eROSITA/EDR page (https://erosita.mpe.mpg.de/edr/eROSITAObservations/Catalogues/. The redshift probability function (PDZs) for each source is also available, under request.
@@ 6 Robustness tests
The middle and right panels of the same Figure shows the performance on the Chandra-COSMOS sample of Marchesi et al. (2016). As expected, the performance for this faint sample is dramatically reduced compared to the X-rays brighter eFEDS sample, given that the counterparts for these fainter X-ray sources will also have faint photometry and thus lower signal-to-noise in the shallow LS10 survey. However, if we limit the sample to sources brighter than about 21 magnitudes (see Table 5), the quality of the photo-z is consistent with previous results, suggesting that we are limited only by the depth of the data.
@@ 6.1 PIT score
For scientific exploitations, in addition to a point prediction, reliable errors must be provided by any photo-z tool. In particular, the full posteriors to reflect our knowledge of the redshift is needed, for example, in the computation of luminosity functions (e.g., Buchner et al., 2015; Fotopoulou et al., 2016; Miyaji et al., 2015). For this reason, we also looked at the Probability Integral Transform (PIT:D’Isanto et al. (2018); Schmidt et al. (2020)) for all three samples. The PIT value is the integral of the posterior up to the spectroscopic redshift (see left panel of Figure 5) and would be uniformly distributed if the spectroscopic redshifts were drawn from the posteriors. The results for the test sample in eFEDS, CSC2, and Chandra-COSMOS are shown in Figure 9. The quantile-quantile plots show the PIT value against a uniformly distributed value. The figures were produced using the Redshift Assessment and Infrastructure Layers (RAIL) framework, discussed in detail in Schmidt et al. (2020). The Figure demonstrates the overall good performance of the posterior distribution and the point prediction performance. It demonstrates that the full posteriors are neither overly nor underly dispersed for eFEDS and CSC2. If the spectroscopic samples were drawn from the posterior, the PIT distribution would be uniform. An excess of PIT values at the lower or upper ends shows that the posteriors are not identifying true values at their lower or upper ends, as it is happening for Chandra-COSMOS. An excess in the center shows that the posteriors are ‘underdispersed’ and underestimate the errors on the point predictions.
@@ 7 Discussion
Thus far, there has been scant optimism regarding the feasibility of attaining accurate photo-zs for AGN in all-sky surveys. This was due to the assumption that many photometric bands from different surveys needed to be assembled and cross-calibrated (unthinkable to be achieved for wide areas). While it remains true that more bands better describe the SED of these complex sources, in this work, we have demonstrated that using an adequate training sample and going beyond the use of fluxes and colors with a machine learning approach, reliable photo-zs can be achieved for the first time, using only optical and WISE data as offered by LS10-South. In particular, the features include information on the light distribution of the sources, parameterized by the fluxes and colors for different apertures, and the quality of the fit assuming different morphological types, in addition to photometric errors and tractorised (Lang, 2014) WISE photometry.
Currently, LS10 is the only survey that provides this kind of information, and we argue here that surveys such as the Legacy Survey of Space and Time (LSST; Ivezić et al., 2019a) and Euclid (Mellier et al., tbs) should provide the same kind of output.
@@ 8 Conclusions and Summary
The model obtained in the training was then applied to the eROSITA/eFEDS field (Section 5). Overall, we have obtained a comparable or better quality photo-z to what was previously available in the best region of the field (REGION 1), with much less effort than what was described in Salvato et al. (2022), despite lacking UV, u and and Near-Infrared band photometry. With the paper, we release the new catalog of photo-zs in eFEDS, including 1 and 3 $\sigma$ errors and PDZ (Section 5.1).
We tested the stability of the results (Section 3.3) by using CircleZ and the model obtained in the training phase on two additional samples, used for a blind test. The first sample, formed by 651 sources from CSC2 in the footprint of LS10-South, shares the same properties (in terms of WISE and X-ray fluxes) as the training sample. There, the quality of the photo-zs is consistent with the one obtained in the validation sample limited to eFEDS sources. The second blind sample consists of all the AGN detected in the Chandra-COSMOS Legacy survey, with reliable spectroscopic redshift and detected in t LS10-South. The metrics indicate that the quality of the photo-zs is consistent with the one computed in the test sample of eFEDS and in CSC2, as long as we consider a limiting AB magnitude of 21.7 in i band. For fainter sources (and thus with larger photometric errors), the accuracy degrades rapidly.
@@ Acknowledgements.
We thank the referee for their careful reading of the manuscript and the constructive feedback. MS thanks Dr. Marianna Lucio for the insights on features analysis and statistics. Part of this work was supported by the German Deutsche Forschungsgemeinschaft, DFG projec Funded by Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2094 – 390783311. CR acknowledges support from Fondecyt Regular grant 1230345 and ANID BASAL project FB210003. RJA was supported by FONDECYT grant number 123171 and the ANID BASAL project FB210003. This work is based on data from eROSITA, the soft X-ray instrument aboard SRG, a joint Russian-German science mission supported by the Russian Space Agency (Roskosmos), in the interests of the Russian Academy of Sciences represented by its Space Research Institute (IKI), and the Deutsches Zentrum für Luft- und Raumfahrt (DLR). The SRG spacecraft was built by Lavochkin Association (NPOL) and its subcontractors and is operated by NPOL with support from the Max Planck Institute for Extraterrestrial Physics (MPE). The development and construction of the eROSITA X-ray instrument were led by MPE, with contributions from the Dr. Karl Remeis Observatory Bamberg & ECAP (FAU Erlangen-Nuernberg), the University of Hamburg Observatory, the Leibniz Institute for Astrophysics Potsdam (AIP), and the Institute for Astronomy and Astrophysics of the University of Tübingen, with the support of DLR and the Max Planck Society. The Argelander Institute for Astronomy of the University of Bonn and the Ludwig Maximilians Universität Munich also participated in the science preparation for eROSITA. The Legacy Surveys consist of three individual and complementary projects: the Dark Energy Camera Legacy Survey (DECaLS; Proposal ID 2014B-0404; PIs: David Schlegel and Arjun Dey), the Beijing-Arizona Sky Survey (BASS; NOAO Prop. ID 2015A-0801; PIs: Zhou Xu and Xiaohui Fan), and the Mayall z-band Legacy Survey (MzLS; Prop. ID 2016A-0453; PI: Arjun Dey). DECaLS, BASS, and MzLS together include data obtained, respectively, at the Blanco telescope, Cerro Tololo Inter-American Observatory, NSF’s NOIRLab; the Bok telescope, Steward Observatory, University of Arizona; and the Mayall telescope, Kitt Peak National Observatory, NOIRLab. Pipeline processing and analyses of the data were supported by NOIRLab and the Lawrence Berkeley National Laboratory (LBNL). The Legacy Surveys project is honored to be permitted to conduct astronomical research on Iolkam Du’ag (Kitt Peak), a mountain with particular significance to the Tohono O’odham Nation. NOIRLab is operated by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation. LBNL is managed by the Regents of the University of California under contract to the U.S. Department of Energy. This project used data obtained with the Dark Energy Camera (DECam), which was constructed by the Dark Energy Survey (DES) collaboration. Funding for the DES Projects has been provided by the U.S. Department of Energy, the U.S. National Science Foundation, the Ministry of Science and Education of Spain, the Science and Technology Facilities Council of the United Kingdom, the Higher Education Funding Council for England, the National Center for Supercomputing Applications at the University of Illinois at Urbana-Champaign, the Kavli Institute of Cosmological Physics at the University of Chicago, Center for Cosmology and Astro-Particle Physics at the Ohio State University, the Mitchell Institute for Fundamental Physics and Astronomy at Texas A & M University, Financiadora de Estudos e Projetos, Fundacao Carlos Chagas Filho de Amparo, Financiadora de Estudos e Projetos, Fundacao Carlos Chagas Filho de Amparo a Pesquisa do Estado do Rio de Janeiro, Conselho Nacional de Desenvolvimento Cientifico e Tecnologico and the Ministerio da Ciencia, Tecnologia e Inovacao, the Deutsche Forschungsgemeinschaft and the Collaborating Institutions in the Dark Energy Survey. The Collaborating Institutions are Argonne National Laboratory, the University of California at Santa Cruz, the University of Cambridge, Centro de Investigaciones Energeticas, Medioambientales y Tecnologicas-Madrid, the University of Chicago, University College London, the DES-Brazil Consortium, the University of Edinburgh, the Eidgenossische Technische Hochschule (ETH) Zurich, Fermi National Accelerator Laboratory, the University of Illinois at Urbana-Champaign, the Institut de Ciencies de l’Espai (IEEC/CSIC), the Institut de Fisica d’Altes Energies, Lawrence Berkeley National Laboratory, the Ludwig Maximilians Universitat Munchen and the associated Excellence Cluster Universe, the University of Michigan, NSF’s NOIRLab, the University of Nottingham, the Ohio State University, the University of Pennsylvania, the University of Portsmouth, SLAC National Accelerator Laboratory, Stanford University, the University of Sussex, and Texas A&M University. BASS is a key project of the Telescope Access Program (TAP), which has been funded by the National Astronomical Observatories of China, the Chinese Academy of Sciences (the Strategic Priority Research Program “The Emergence of Cosmological Structures” Grant # XDB09000000), and the Special Fund for Astronomy from the Ministry of Finance. The BASS is also supported by the External Cooperation Program of Chinese Academy of Sciences (Grant # 114A11KYSB20160057), and Chinese National Natural Science Foundation (Grant # 12120101003, # 11433005). The Legacy Survey team uses data products from the Near-Earth Object Wide-field Infrared Survey Explorer (NEOWISE), a project of the Jet Propulsion Laboratory/California Institute of Technology. NEOWISE is funded by the National Aeronautics and Space Administration. The Legacy Surveys imaging of the DESI footprint is supported by the Director, Office of Science, Office of High Energy Physics of the U.S. Department of Energy under Contract No. DE-AC02-05CH1123, by the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility under the same contract, and by the U.S. National Science Foundation, Division of Astronomical Sciences under Contract No. AST-0950945 to NOAO. Funding for the Sloan Digital Sky Survey V has been provided by the Alfred P. Sloan Foundation, the Heising-Simons Foundation, the National Science Foundation, and the Participating Institutions. SDSS acknowledges support and resources from the Center for High-Performance Computing at the University of Utah. SDSS telescopes are located at Apache Point Observatory, funded by the Astrophysical Research Consortium and operated by New Mexico State University, and at Las Campanas Observatory, operated by the Carnegie Institution for Science. The SDSS web site is www.sdss.org. SDSS is managed by the Astrophysical Research Consortium for the Participating Institutions of the SDSS Collaboration, including Caltech, The Carnegie Institution for Science, Chilean National Time Allocation Committee (CNTAC) ratified researchers, The Flatiron Institute, the Gotham Participation Group, Harvard University, Heidelberg University, The Johns Hopkins University, L’Ecole polytechnique fédérale de Lausanne (EPFL), Leibniz-Institut für Astrophysik Potsdam (AIP), Max-Planck-Institut für Astronomie (MPIA Heidelberg), Max-Planck-Institut für Extraterrestrische Physik (MPE), Nanjing University, National Astronomical Observatories of China (NAOC), New Mexico State University, The Ohio State University, Pennsylvania State University, Smithsonian Astrophysical Observatory, Space Telescope Science Institute (STScI), the Stellar Astrophysics Participation Group, Universidad Nacional Autónoma de México, University of Arizona, University of Colorado Boulder, University of Illinois at Urbana-Champaign, University of Toronto, University of Utah, University of Virginia, Yale University, and Yunnan University.

### v2

@@ 3.2 Extension of the training sample to high-z sources
The contribution to the high-redshift AGN population at $z$¿5 from X-ray pencil-beam surveys is feeble  [^16] Unsurprisingly, most of the sources at redshift higher than five that we visually inspected were wrong. By construction, the training sample would not represent the high-$z$ population. However, eROSITA is expected to detect about 90 sources at $z$¿5.5 by the end of the mission (Wolf et al., 2021). At the same time, about a dozen have already been identified (Medvedev et al., 2020, 2021; Wolf et al., 2021, 2023, 2024). The optical spectra of these high-z QSOs closely resemble the spectra of QSOs at low redshift (e.g., Mortlock et al., 2011; Yang et al., 2021), which is why we have extended the training sample, adding 283 sources spectroscopically confirmed at $z>5.5$, from the compilation of Fan et al. (2023), regardless of whether or not they are detected in the X-rays. The violet dashed histogram in Figure 3 shows the redshift distribution of these sources.
Table 2: Sources used in this work
@@
 [^17] The table shows the breakdown distribution of sources used for training (i.e., training, test, and validation samples) and in the blind samples, before and after the match with the spectroscopic redshift compilation, and after cleaning for duplicates. The numbers within the parenthesis show the contribution to the sample used for training.
@@ 3.3 Samples for blind tests
CSC2: this blind sample is generated using the counterparts to the CSC2 objects [^18] https://cxc.cfa.harvard.edu/csc/ in the same way as described for the training sample. The original sample of CSC2 sources with reliable counterparts included 5643 sources. Screening for sources with spectroscopic redshift and after removing the sources that are in common with the training sample, 413 remain in the area covered by LS10-South, after including 26 redshifts from SDSS-V.
@@
Figure 4: X-rays flux vs. W1 magnitude distributions for the training (left panel), and the blind samples: CSC2 (middle panel) and Chandra-COSMOS (right panel). The dashed line represents the galactic/extragalactic separation proposed in (Salvato et al., 2018b), who already pointed out that bright nearby galaxies would be misclassified. The contours encompassing 50%, 80%, and 95% of the training sample distribution are over-plotted on all panels. While the CSC2 sample has a distribution similar to the training sample, the distribution of the Chandra-COSMOS Legacy sources is biased toward fainter values of X-rays and W1.
@@ 4.1 Problem Formulation
However, as shown in D’Isanto et al. (2018), a single trained network producing a PDF cannot capture the uncertainty. We, therefore, extended our approach for uncertainty estimation and followed Lakshminarayanan et al. (2017) who proposed to use an ensemble of deep neural networks such as MDNs to improve uncertainty estimates. We follow this approach with an ensemble of MDNs that, given the photometric input features, each predicts the parameters of the Gaussian mixture distribution with a certain number of components. To be specific, the $i^{th}$ model $f_{i}(x;\theta_{i})$ in an ensemble of $N$ models, given its parameters $\theta_{i}$ and the input features $x$, predicts the parameters of a Gaussian mixture composed of $n_{i}$ Gaussian, defined by the means $\mu^{i}(x;\theta_{i})$, and standard deviations $\sigma^{i}(x;\theta_{i})$ and weights $\omega^{i}(x;\theta_{i})$. The conditional predictive distribution defined by this MDN is given by:
@@ 4.2.2 Training procedure and hyperparameter search
Table 3: Model and training hyperparameters with the range within which they are selected.
@@
 [^20] These specific ranges are selected by observing model performance with hyperparameter values in wider ranges.
@@ 4.2.3 Inference and post-processing
Table 4: Averaged statistics across all individually trained NLL loss and CPRS models.
@@
222222the table includes aggregated information for the combined ensemble of all NLL and CRPS models. Furthermore, the fraction of outliers and accuracy are listed for the best-performing ensembles.
@@ 5 Redshift computation for eFEDS sources and comparison with literature
It is also worth mentioning that the metrics reported for LePHARE and DNNZ were measured using the training sample, not on the test and/or validation sample. If we compute the metrics for CircleZ using all the spectroscopy, i.e., including those used for the training and validation, the values would suggest a quality close to the one available for the best pencil-beam surveys (XMM-COSMOS, (E)CDF Salvato et al., 2009; Luo et al., 2010; Hsu et al., 2014), computed with deep, rich and homogenized photometry). This assures that the new photo-zs made available for eFEDS in this paper are reliable and can be used for Luminosity Function studies or for preselecting specific instruments or wavelengths of interest when submitting a proposal.
@@
Table 5: Comparison of bias, fraction of outliers, and accuracy between photo-z computed with CircleZ, LePHARE, and DNNZ, for extragalactic sources in the eROSITA/eFEDS field.
@@
 [^23] The values for LePHARE and DNNZ are taken directly from (Salvato et al., 2022) for comparison. While the selection of the templates for LePHARE is based on the same sample used for assessing the quality of photo-zs, the training for DNNz(CircleZ) is partially(completely) independent from the test and validation sample used here. The quality of CircleZ exceeds the one obtained by both LePHARE and DNNZ in the area outside KiDS, demonstrating the strength of the method in general and particularly where there is a lack of near-infrared photometry.
@@
sectionNew photo-z catalog release for eFEDS With this paper, we release a new catalog of photo-z computed with CircleZ, for the extragalactic sources in the eROSITA/eFEDS field. For each source we provide, in addition to the eROSITA_ID, the photo-z, and the 1 and 3 $\sigma$ lower and upper errors, also a unique identifier for the counterpart in the LS10 catalog. The identifier is created by concatenating the columns RELEASE, BRICKID, and OBJID. In addition, we provide about 5000 new spectroscopic redshifts that became available with the 18th Data Release of SDSS-V (Almeida et al., 2023). The catalog is available via Vizier and on the eROSITA/EDR page [^24] https://erosita.mpe.mpg.de/edr/eROSITAObservations/Catalogues/. The redshift probability function (PDZs) for each source is also available, under request.
@@ 6 Robustness tests
The middle and right panels of the same Figure show the performance on the Chandra-COSMOS sample of Marchesi et al. (2016). As expected, the performance for this faint sample is dramatically reduced compared to the X-rays brighter eFEDS sample, given that the counterparts for these fainter X-ray sources will also have faint photometry and thus lower signal-to-noise in the shallow LS10 survey. However, if we limit the sample to sources brighter than about 21 magnitudes (see Table 23), the quality of the photo-z is consistent with previous results, suggesting that we are limited only by the depth of the data.
@@ 6.1 PIT score
For scientific exploitations, in addition to a point prediction, reliable errors must be provided by any photo-z tool. In particular, the full posteriors to reflect our knowledge of the redshift is needed, for example, in the computation of luminosity functions (e.g., Buchner et al., 2015; Fotopoulou et al., 2016; Miyaji et al., 2015). For this reason, we also looked at the Probability Integral Transform (PIT:D’Isanto et al. (2018); Schmidt et al. (2020)) for all three samples. The PIT value is the integral of the posterior up to the spectroscopic redshift (see left panel of Figure 5) and would be uniformly distributed if the spectroscopic redshifts were drawn from the posteriors. The results for the test sample in eFEDS, CSC2, and Chandra-COSMOS are shown in Figure 9. The quantile-quantile plots show the PIT value against a uniformly distributed value. The figures were produced using the Redshift Assessment and Infrastructure Layers (RAIL) framework, discussed in detail in Schmidt et al. (2020). The Figure demonstrates the overall good performance of the posterior distribution and the point prediction performance. It demonstrates that the full posteriors are neither overly nor underly dispersed for eFEDS and CSC2. If the spectroscopic samples were drawn from the posterior, the PIT distribution would be uniform. An excess of PIT values at the lower or upper ends shows that the posteriors are not identifying true values at their lower or upper ends, as it is happening for Chandra-COSMOS. An excess in the center shows that the posteriors are underdispersed and underestimate the errors on the point predictions.
@@ 7 Discussion
Thus far, there has been scant optimism regarding the feasibility of attaining accurate photo-zs for AGN in all-sky surveys. This was due to the assumption that many photometric bands from different surveys needed to be assembled and cross-calibrated (unthinkable to be achieved for wide areas). While it remains true that more bands better describe the SED of these complex sources, in this work, we have demonstrated that using an adequate training sample and going beyond the use of fluxes and colors with a machine learning approach, reliable photo-zs can be achieved for the first time, using only optical and WISE data as offered by LS10-South. In particular, the features include information on the light distribution of the sources, parameterized by the fluxes and colors for different apertures, and the quality of the fit assuming different morphological types, in addition to photometric errors and tractorized (Lang, 2014) WISE photometry.
Currently, LS10 is the only survey that provides this kind of information, and we argue here that surveys such as the Legacy Survey of Space and Time (LSST; Ivezić et al., 2019a) and Euclid (Euclid Collaboration et al., 2024) should provide the same kind of output.
@@ 8 Conclusions and Summary
The model obtained in the training was then applied to the eROSITA/eFEDS field (Section 5). Overall, we have obtained a comparable or better quality photo-z to what was previously available in the best region of the field (REGION 1), with much less effort than what was described in Salvato et al. (2022), despite lacking UV, u and and Near-Infrared band photometry. With the paper, we release the new catalog of photo-zs in eFEDS, including 1 and 3 $\sigma$ errors and PDZ (Section 7).
We tested the stability of the results (Section 3.3) by using CircleZ and the model obtained in the training phase on two additional samples, used for a blind test. The first sample, formed by 651 sources from CSC2 in the footprint of LS10-South, shares the same properties (in terms of WISE and X-ray fluxes) as the training sample. There, the quality of the photo-zs is consistent with the one obtained in the validation sample limited to eFEDS sources. The second blind sample consists of all the AGN detected in the Chandra-COSMOS Legacy survey, with reliable spectroscopic redshift and detected in t LS10-South. The metrics indicate that the quality of the photo-zs is consistent with the one computed in the test sample of eFEDS and in CSC2, as long as we consider a limiting AB magnitude of 21.7 in i band. The accuracy degrades rapidly for fainter sources (and thus with larger photometric errors).
@@ Acknowledgements.
We thank the referee for their careful reading of the manuscript and the constructive feedback. MS thanks Dr. Marianna Lucio for the insights on features analysis and statistics. Part of this work was supported by the German Deutsche Forschungsgemeinschaft, DFG project Funded by Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2094 – 390783311. CR acknowledges support from Fondecyt Regular grant 1230345 and ANID BASAL project FB210003. RJA was supported by FONDECYT grant number 1231718 and the ANID BASAL project FB210003. This work is based on data from eROSITA, the soft X-ray instrument aboard SRG, a joint Russian-German science mission supported by the Russian Space Agency (Roskosmos), in the interests of the Russian Academy of Sciences represented by its Space Research Institute (IKI), and the Deutsches Zentrum für Luft- und Raumfahrt (DLR). The SRG spacecraft was built by Lavochkin Association (NPOL) and its subcontractors and is operated by NPOL with support from the Max Planck Institute for Extraterrestrial Physics (MPE). The development and construction of the eROSITA X-ray instrument were led by MPE, with contributions from the Dr. Karl Remeis Observatory Bamberg & ECAP (FAU Erlangen-Nuernberg), the University of Hamburg Observatory, the Leibniz Institute for Astrophysics Potsdam (AIP), and the Institute for Astronomy and Astrophysics of the University of Tübingen, with the support of DLR and the Max Planck Society. The Argelander Institute for Astronomy of the University of Bonn and the Ludwig Maximilians Universität Munich also participated in the science preparation for eROSITA. The Legacy Surveys consist of three individual and complementary projects: the Dark Energy Camera Legacy Survey (DECaLS; Proposal ID 2014B-0404; PIs: David Schlegel and Arjun Dey), the Beijing-Arizona Sky Survey (BASS; NOAO Prop. ID 2015A-0801; PIs: Zhou Xu and Xiaohui Fan), and the Mayall z-band Legacy Survey (MzLS; Prop. ID 2016A-0453; PI: Arjun Dey). DECaLS, BASS, and MzLS together include data obtained, respectively, at the Blanco telescope, Cerro Tololo Inter-American Observatory, NSF’s NOIRLab; the Bok telescope, Steward Observatory, University of Arizona; and the Mayall telescope, Kitt Peak National Observatory, NOIRLab. Pipeline processing and analyses of the data were supported by NOIRLab and the Lawrence Berkeley National Laboratory (LBNL). The Legacy Surveys project is honored to be permitted to conduct astronomical research on Iolkam Du’ag (Kitt Peak), a mountain with particular significance to the Tohono O’odham Nation. NOIRLab is operated by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation. LBNL is managed by the Regents of the University of California under contract to the U.S. Department of Energy. This project used data obtained with the Dark Energy Camera (DECam), which was constructed by the Dark Energy Survey (DES) collaboration. Funding for the DES Projects has been provided by the U.S. Department of Energy, the U.S. National Science Foundation, the Ministry of Science and Education of Spain, the Science and Technology Facilities Council of the United Kingdom, the Higher Education Funding Council for England, the National Center for Supercomputing Applications at the University of Illinois at Urbana-Champaign, the Kavli Institute of Cosmological Physics at the University of Chicago, Center for Cosmology and Astro-Particle Physics at the Ohio State University, the Mitchell Institute for Fundamental Physics and Astronomy at Texas A & M University, Financiadora de Estudos e Projetos, Fundacao Carlos Chagas Filho de Amparo, Financiadora de Estudos e Projetos, Fundacao Carlos Chagas Filho de Amparo a Pesquisa do Estado do Rio de Janeiro, Conselho Nacional de Desenvolvimento Cientifico e Tecnologico and the Ministerio da Ciencia, Tecnologia e Inovacao, the Deutsche Forschungsgemeinschaft and the Collaborating Institutions in the Dark Energy Survey. The Collaborating Institutions are Argonne National Laboratory, the University of California at Santa Cruz, the University of Cambridge, Centro de Investigaciones Energeticas, Medioambientales y Tecnologicas-Madrid, the University of Chicago, University College London, the DES-Brazil Consortium, the University of Edinburgh, the Eidgenossische Technische Hochschule (ETH) Zurich, Fermi National Accelerator Laboratory, the University of Illinois at Urbana-Champaign, the Institut de Ciencies de l’Espai (IEEC/CSIC), the Institut de Fisica d’Altes Energies, Lawrence Berkeley National Laboratory, the Ludwig Maximilians Universitat Munchen and the associated Excellence Cluster Universe, the University of Michigan, NSF’s NOIRLab, the University of Nottingham, the Ohio State University, the University of Pennsylvania, the University of Portsmouth, SLAC National Accelerator Laboratory, Stanford University, the University of Sussex, and Texas A&M University. BASS is a key project of the Telescope Access Program (TAP), which has been funded by the National Astronomical Observatories of China, the Chinese Academy of Sciences (the Strategic Priority Research Program “The Emergence of Cosmological Structures” Grant # XDB09000000), and the Special Fund for Astronomy from the Ministry of Finance. The BASS is also supported by the External Cooperation Program of Chinese Academy of Sciences (Grant # 114A11KYSB20160057), and Chinese National Natural Science Foundation (Grant # 12120101003, # 11433005). The Legacy Survey team uses data products from the Near-Earth Object Wide-field Infrared Survey Explorer (NEOWISE), a project of the Jet Propulsion Laboratory/California Institute of Technology. NEOWISE is funded by the National Aeronautics and Space Administration. The Legacy Surveys imaging of the DESI footprint is supported by the Director, Office of Science, Office of High Energy Physics of the U.S. Department of Energy under Contract No. DE-AC02-05CH1123, by the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility under the same contract, and by the U.S. National Science Foundation, Division of Astronomical Sciences under Contract No. AST-0950945 to NOAO. Funding for the Sloan Digital Sky Survey V 272727www.sdss.org has been provided by the Alfred P. Sloan Foundation, the Heising-Simons Foundation, the National Science Foundation, and the Participating Institutions. SDSS acknowledges support and resources from the Center for High-Performance Computing at the University of Utah. SDSS telescopes are located at Apache Point Observatory, funded by the Astrophysical Research Consortium and operated by New Mexico State University, and at Las Campanas Observatory, operated by the Carnegie Institution for Science. SDSS is managed by the Astrophysical Research Consortium for the Participating Institutions of the SDSS Collaboration, including Caltech, The Carnegie Institution for Science, Chilean National Time Allocation Committee (CNTAC) ratified researchers, The Flatiron Institute, the Gotham Participation Group, Harvard University, Heidelberg University, The Johns Hopkins University, L’Ecole polytechnique fédérale de Lausanne (EPFL), Leibniz-Institut für Astrophysik Potsdam (AIP), Max-Planck-Institut für Astronomie (MPIA Heidelberg), Max-Planck-Institut für Extraterrestrische Physik (MPE), Nanjing University, National Astronomical Observatories of China (NAOC), New Mexico State University, The Ohio State University, Pennsylvania State University, Smithsonian Astrophysical Observatory, Space Telescope Science Institute (STScI), the Stellar Astrophysics Participation Group, Universidad Nacional Autónoma de México, University of Arizona, University of Colorado Boulder, University of Illinois at Urbana-Champaign, University of Toronto, University of Utah, University of Virginia, Yale University, and Yunnan University.